### v1

# Generative Ontology: When Structured Knowledge Learns to Create

Abstract
Traditional ontologies excel at describing domain structure but cannot generate novel artifacts. Large language models generate fluently but produce outputs that lack structural validity, hallucinating mechanisms without components, goals without end conditions. We introduce Generative Ontology, a framework that synthesizes these complementary strengths: ontology provides the grammar; the LLM provides the creativity.
Generative Ontology encodes domain knowledge as executable Pydantic schemas that constrain LLM generation via DSPy signatures. A multi-agent pipeline assigns specialized roles to different ontology domains: a Mechanics Architect designs game systems, a Theme Weaver integrates narrative, a Balance Critic identifies exploits. Each agent carrying a professional “anxiety” that prevents shallow, agreeable outputs. Retrieval-augmented generation grounds novel designs in precedents from existing exemplars, while iterative validation ensures coherence between mechanisms and components.
We demonstrate the framework through GameGrammar, a system for generating complete tabletop game designs. Given a thematic prompt (“bioluminescent fungi competing in a cave ecosystem”), the pipeline produces structurally complete, playable game specifications with mechanisms, components, victory conditions, and setup instructions. These outputs satisfy ontological constraints while remaining genuinely creative.
The pattern generalizes beyond games. Any domain with expert vocabulary, validity constraints, and accumulated exemplars (music composition, software architecture, culinary arts) is a candidate for Generative Ontology. We argue that constraints do not limit creativity but enable it: just as grammar makes poetry possible, ontology makes structured generation possible.
1  Introduction
Ontologies describe what is. They define concepts, relationships, and constraints that codify expert understanding of a domain. In our earlier work on tabletop game ontology [5], we built a structured vocabulary that could decompose Catan into mechanisms (resource trading, modular board, dice-driven production), components (hex tiles, resource cards, settlements), and player dynamics (competitive, negotiation-heavy, variable player count). That ontology gave us a precise language for analysis.
But analysis is not creation.
Understanding the structure of sonnets does not make one a poet. Knowing the rules of chess does not generate new games. A traditional ontology, for all its precision, remains passive, a tool for describing what already exists, not for imagining what could be. We ask: what happens when we transform that passive vocabulary into an active grammar?
1.1  The Problem: Structural Hallucination
Large language models (LLMs) excel at fluent generation. Ask an LLM to “design a deck-building game set in a haunted mansion” and it will produce paragraphs of creative output: themes, mechanisms, components, and victory conditions flowing freely. Yet this fluency hides a deeper problem: LLMs generate without understanding structure.
Consider the output for such a prompt: “Players explore Ravenshollow Manor, collecting ghost cards and building their decks. The game features a unique ‘fear mechanic’ where players manage their sanity. Victory goes to whoever escapes with the most treasure…” This sounds plausible. But what cards exist in the starting deck? How do players acquire new cards? What triggers the end of the game? How does the “fear mechanic” interact with deck composition? The LLM has generated the appearance of a game design without the substance. Handed to a playtester, it would collapse into confusion. This is the hallucination problem applied to creative domains: just as LLMs hallucinate facts, they hallucinate structures, producing outputs that sound valid without being valid.
Figure 1: A sample Game Ontology expressed as an OntoUML class diagram. The Game root entity decomposes into Player, Mechanism, and Component aggregates, each with typed relationships (hasPlayer, hasMechanism, hasComponent) that mirror the domain structure a Generative Ontology must capture and enforce.
1.2  Complementary Limitations
Table 1 summarizes the impasse. Traditional ontology has structure without creativity; the LLM has creativity without structure. Neither alone solves the problem of generating novel, valid designs.
Table 1: Traditional ontology versus pure LLM generation.
Strength Weakness
Ontology Precise, structured, validated Cannot generate novel outputs
LLM Creative, fluent, abundant Unstructured, hallucinated
The insight of Generative Ontology is that these limitations are complementary. What ontology lacks, LLMs provide. What LLMs lack, ontology provides. The synthesis, structured generation guided by ontological schemas, produces outputs that are both creative and valid.
1.3  Related Work
Ontology development has a well-established methodology [14], but recent research at the intersection of ontologies and LLMs proceeds primarily in the opposite direction from our approach. DRAGON-AI uses retrieval-augmented LLM generation to construct biomedical ontologies from text corpora [15], and OLLM demonstrates end-to-end ontology learning with language models. These systems use LLMs to generate ontologies; we use ontologies to guide generation.
Mehenni and Zouaq [12] demonstrate ontology-constrained generation for clinical summaries, the closest methodological precedent to our work. However, their goal is hallucination reduction in summarization, whereas we target structured creativity. In game generation, Gallotta et al. [9] survey LLM applications in games broadly, while Becker et al. [4] present Boardwalk, a framework for generating board game code from existing rules, generating implementations, not designs. Maleki and Zhao [11] survey procedural content generation methods, noting emerging LLM integration but without ontological constraints. Cooper [7] provides foundational work on constrained decoding that underpins our structured output approach.
Our contribution fills a specific gap: using ontology as a generative grammar for structured creative output, operationalized through executable schemas, multi-agent pipelines, and retrieval augmentation.
1.4  Paper Overview
We introduce Generative Ontology and develop it through the domain of tabletop game design. Section 2 defines the framework and its philosophical grounding. Section 3 transforms ontology classes into executable Pydantic schemas. Section 4 operationalizes generation with DSPy signatures. Section 5 introduces multi-agent specialization with professional anxieties. Section 6 adds retrieval-augmented grounding. Section 7 develops the ontology as a validation contract. Section 8 presents a complete case study. Section 9 argues the pattern generalizes beyond games. Section 10 concludes with future directions.
2  Generative Ontology: The Synthesis
If traditional ontology is a map, Generative Ontology is a map that knows how to build new cities.
Generative Ontology is the practice of encoding domain knowledge as executable schemas that constrain and guide AI generation, transforming static knowledge representation into a creative engine.
When we built our tabletop game ontology [5], we created a structured vocabulary: game types, mechanisms, components, player interactions. This vocabulary enabled analysis of games such as Catan and Dune: Imperium with precision. The insight of Generative Ontology is that the same structure enabling understanding can enable creation. The categories that constrain analysis become the scaffolding for generation.
2.1  The Grammar of Games
A poet does not experience grammar as a limitation. Grammar is not what prevents poetry; it is what makes poetry possible. Without the structure of syntax, semantics, and form, there would be no sonnets, no haiku, no free verse pushing against convention. The rules are not the enemy of creativity; they are its condition of possibility.
The same principle applies to game design. When we ask an unconstrained LLM to “design a board game about space exploration,” we receive verbose descriptions that sound plausible but collapse under scrutiny. The AI might describe a “resource management mechanic” without specifying what resources exist, how they flow, or what makes acquiring them interesting. Generative Ontology provides the grammar. When we encode the game ontology as a schema, we give the LLM the structural vocabulary to be creative coherently. The schema declares: every game must have a goal, an end condition, mechanisms that create player choices, components that instantiate those mechanisms. Within those constraints, infinitely many games are possible. Without them, no valid game emerges.
Figure 2: LLM potential plus ontology constraints yields valid design. Generative Ontology transforms a passive vocabulary for description into an active grammar for creation.
Table 2 illustrates the difference concretely.
Table 2: Generation without versus with Generative Ontology.
Without Ontology With Generative Ontology
“Players collect resources and build things” Goal: first to 10 VP via settlements
“There’s some kind of trading” Mechanism: resource trading with 4:1 bank or negotiated rates
“The game ends eventually” End Condition: when any player reaches 10 VP or development deck exhausts
Fluent but vague Structured and playable
2.2  Philosophical Grounding
For readers familiar with our earlier exploration of process philosophy for AI agent design [6], Generative Ontology has a natural grounding in Alfred North Whitehead’s metaphysics [16]. Whitehead distinguished between eternal objects (pure forms, patterns, and structures that exist as potentials) and actual occasions (the concrete events where eternal objects find expression). Our game ontology is a collection of eternal objects: the abstract patterns of worker placement, deck building, area control. These patterns have no reality until they are instantiated in actual games: in Catan’s modular hex board, in Dune: Imperium’s faction tracks. Generative Ontology is the process of moving from eternal objects to actual occasions: the ontology provides the forms; generation produces the instances. Creativity, in Whitehead’s view, is precisely this novel synthesis of eternal objects into actual occasions that have never existed before [3]. The forms constrain but do not determine. Each generated game is a genuine novelty.
3  From Ontology to Executable Schema
Philosophy illuminates the path; engineering builds the road. The key insight is surprisingly direct: ontology classes map naturally to schema definitions, and schema definitions become the structured outputs that constrain LLM generation.
3.1  The Transformation
Our tabletop game ontology [5, 8] defined four core concepts: Game (the container with goal and end condition), Mechanism (turn structure, actions, resolution, uncertainty), Component (board, cards, tokens), and Player (roles and interactions). Each of these ontology classes becomes a Pydantic BaseModel, a typed, validated structure that LLMs must conform to when generating output.
Listing 1 shows the abbreviated GameOntology schema. The MechanismType enumeration reflects ontological categories developed over decades of game design taxonomy [8]. By encoding these as an enum, we ensure the LLM can only reference mechanisms within our shared vocabulary. Every field in GameOntology corresponds to an ontology concept. The goal and end_condition are required strings with minimum lengths, preventing vague outputs. The nested ComponentSet and PlayerDynamics models enforce structure at every level.
1from enum import Enum
2from typing import List, Literal, Optional
3from pydantic import BaseModel, Field
4
5class MechanismType(str, Enum):
6 WORKER_PLACEMENT = "worker_placement"
7 ACTION_POINTS = "action_points"
8 DECK_BUILDING = "deck_building"
9 RESOURCE_MGMT = "resource_management"
10 AREA_CONTROL = "area_control"
11 ENGINE_BUILDING = "engine_building"
12 SET_COLLECTION = "set_collection"
13 HIDDEN_INFO = "hidden_information"
14
15class GameOntology(BaseModel):
16 title: str = Field(description="Name of the game")
17 theme: str = Field(description="Setting and narrative")
18 game_type: Literal["cooperative","competitive",
19 "semi-cooperative"]
20 goal: str = Field(min_length=20,
21 description="Victory condition")
22 end_condition: str = Field(min_length=10,
23 description="What triggers game end")
24 primary_mechanisms: List[MechanismType] = Field(
25 min_length=2, max_length=4)
26 turn_structure: str
27 uncertainty_source: str
28 components: ComponentSet # nested model
29 players: PlayerDynamics # nested model
30 setup: str
31 core_loop: str
32 strategic_depth: str
Listing 1: Abbreviated GameOntology Pydantic schema. Each field corresponds to an ontology concept; enums constrain the LLM to recognized mechanisms.
3.2  Ontology-to-Schema Correspondence
Table 3 formalizes the mapping. The correspondence is not accidental: every ontology class finds expression in the schema, and every ontological relationship (Game has Mechanisms, Game uses Components) becomes a nested field or list. The schema is the ontology made executable.
Table 3: Ontology-to-schema correspondence.
Concept Schema Type Purpose
Game Types Literal[...] Valid modes
Mechanisms List[Mechanism] Mechanics
Components ComponentSet Physical parts
Goal / End str (required) Playability
Players PlayerDynamics Interactions
3.3  Why Pydantic?
The choice of Pydantic [2] is deliberate. Declaring GameOntology as a BaseModel provides four critical capabilities: (1) type validation, where the LLM’s output is automatically validated against the schema; (2) self-documentation, where each Field description becomes part of the prompt context; (3) nested structure, where hierarchical ontologies map naturally to nested models; and (4) enum constraints, where the LLM operates within the ontological vocabulary and cannot invent mechanisms outside the taxonomy.
The original ontology was a tool for analysis: we could take Catan and decompose it into ontology terms. The schema flips this relationship. Now the ontology becomes a tool for synthesis: we provide a theme, and the constrained generation produces a complete game that satisfies every ontological requirement.
4  Operationalizing with DSPy
We have the ontology encoded as Pydantic schemas. Now we need an engine to drive generation. DSPy (Declarative Self-improving Python) [10] treats LLM interactions as composable, typed operations rather than monolithic prompt templates. This philosophy aligns with Generative Ontology: DSPy signatures are ontological contracts.
4.1  Signatures as Ontological Contracts
A DSPy signature declares inputs, outputs, and a docstring that establishes the LLM’s persona. Listing 2 shows the core generation signature. The docstring becomes the system prompt; the input fields capture what the user provides; the output field declares that the result must conform to our GameOntology schema.
1import dspy
2from game_ontology import GameOntology
3
4class GenerativeOntologySignature(dspy.Signature):
5 """You are an expert tabletop game designer
6 with deep knowledge of mechanisms, player
7 dynamics, and component design. Design a
8 complete, playable game based on the theme.
9 Every field must be specific, not vague."""
10
11 theme = dspy.InputField(
12 desc="Thematic concept for the game")
13 design_constraints = dspy.InputField(
14 desc="Player count, complexity, time",
15 default="")
16 game_design: GameOntology = dspy.OutputField(
17 desc="Complete game design per schema")
Listing 2: DSPy GenerativeOntologySignature. The typed output field binds generation to the ontological schema.
4.2  Execution Flow
When we execute this signature, DSPy performs the following sequence: (1) constructs a prompt incorporating the docstring, field descriptions, and the full Pydantic schema structure; (2) sends the prompt to the configured LLM; (3) parses the response into the GameOntology model; (4) validates that all required fields are present and correctly typed; and (5) returns the validated result or raises an error for retry.
DSPy’s ChainOfThought module extends this by instructing the model to reason through the design before committing to structured output. The reasoning captures why the system chose worker placement over deck building, how the theme informed the end condition, and what creates tension in the player dynamics. This reasoning is not merely debugging output; it documents the design rationale.
Encapsulating logic in a DSPy Module enables composition of multiple generation steps, post-generation validation, and retry logic. The module adds a layer of semantic validation beyond what Pydantic provides: the schema ensures structural correctness (all fields present, correct types), while the module ensures coherence: that mechanisms match components and that game type aligns with player dynamics. This layered architecture (schema for structure, module for semantics) is central to the reliability of Generative Ontology.
5  Multi-Agent Ontology Pipeline
A single LLM call, even with our ontology schema, carries the entire burden of game design: it must simultaneously consider mechanisms, theme, components, balance, and player experience. Human design teams do not work this way. A designer sketches mechanics while an artist develops visual language while a playtester identifies broken interactions. Generative Ontology enables the same division of labor by decomposing the ontology into domains and assigning specialized agents to each.
5.1  Agent Roster and Professional Anxieties
Table 4 lists the five agents. The “Anxiety” column is crucial: each agent is not merely assigned a domain but is given a concern, a professional worry that shapes its generation and critique. The Mechanics Architect does not just produce mechanisms; it worries about whether those mechanisms create meaningful choices. This anxiety-driven design prevents the “yes-man” tendency of LLMs to produce plausible but shallow output.
Table 4: Agent roster with ontology domains and professional anxieties.
Agent Domain Anxiety
Mechanics Architect Mechanisms, turn structure “Is there meaningful player agency?”
Theme Weaver Narrative, setting “Does the theme feel alive in every mechanism?”
Component Designer Cards, tokens, board “Can players manipulate this smoothly?”
Balance Critic Cross-domain analysis “What breaks when optimized?”
Fun Factor Judge Player experience “Would I want to play this again?”
Each agent receives a DSPy signature tailored to its domain, with the docstring encoding its persona and anxiety. The Mechanics Architect’s signature begins: “You are a veteran game mechanics designer who has studied hundreds of tabletop games. You obsess over player agency…” The Balance Critic’s opens: “You are a competitive player and game breaker. Your job is to find what’s broken…” These personas are not decorative; they shape the distribution of generated outputs toward domain-appropriate rigor.
5.2  Pipeline Architecture
Figure 3: Multi-agent pipeline architecture. Generative agents build sequentially (Mechanics $\to$ Theme $\to$ Components); critic agents evaluate in parallel; refinement addresses identified issues before final assessment.
The pipeline, illustrated in Figure 3, proceeds in four phases:
  1. 1.
    Sequential Generation. The Mechanics Architect produces core systems given the theme and constraints. The Theme Weaver receives these mechanics and integrates narrative. The Component Designer receives both and specifies physical elements. Each agent builds on the previous, with information flowing through typed Pydantic schemas.
  2. 2.
    Critical Evaluation. The Balance Critic examines the assembled design for exploits, dominant strategies, and interaction gaps. Issues are tagged with severity levels (minor, moderate, severe).
  3. 3.
    Refinement. If moderate or severe issues are found, a refinement agent addresses specific recommendations while preserving the design’s strengths.
  4. 4.
    Experience Assessment. The Fun Factor Judge evaluates engagement hooks, tension moments, satisfaction sources, and replayability, producing a 1–10 rating and improvement suggestions.
5.3  Why Multi-Agent Outperforms Single-Pass
The multi-agent approach offers five advantages: (1) focused expertise, where each agent optimizes for its domain without distraction; (2) built-in tension, where agents have different anxieties, mirroring real design team dynamics; (3) explicit handoffs, where information flows through defined schema interfaces; (4) iterative refinement, where critics identify issues that refiners address; and (5) observability, where when something fails, we know which agent to adjust.
6  Retrieval-Augmented Generation
Our Generative Ontology can create novel game designs, but creativity without context risks reinventing solutions that already exist, or repeating mistakes that experienced designers have learned to avoid. The most interesting creative work happens not in a vacuum but in conversation with what has come before.
Figure 4: Knowledge System architecture. A unified Retriever Service dispatches queries to ChromaDB (semantic vector search) or SQLite (structured fallback), drawing on a BGG corpus of 2,000 games with 384-dimensional embeddings, mechanism types, and filterable metadata.
6.1  Two-Phase Retrieval
BoardGameGeek [1] contains data on over 100,000 tabletop games. We structure this corpus according to our ontology and employ a two-phase retrieval strategy: (1) ontology filtering, which narrows candidates to games sharing target mechanisms from the MechanismType taxonomy; then (2) semantic ranking, which ranks the filtered set by embedding similarity to the thematic query. This ensures retrieved games are both mechanically relevant and thematically resonant.
As a concrete example, for the prompt “trading empires in the ancient Mediterranean” with target mechanisms [TRADING, AREA_CONTROL, RESOURCE_MANAGEMENT], the system retrieves Concordia (card-driven trading networks across the Roman Empire), Cyclades (Greek island control with auction-based resource competition), and Tigris & Euphrates (civilization building with multi-dimensional area control scoring). The generated design synthesizes insights from Concordia’s card mechanism and Tigris & Euphrates’ balanced scoring into something novel that explicitly acknowledges its lineage in a design rationale field.
6.2  Grounded Creativity
RAG transforms Generative Ontology from invention to informed invention. The system knows what mechanisms have been successfully combined, how similar themes have been approached, and what design patterns work for given player counts and complexity levels. When retrieval returns few results for a mechanism–theme combination, we have discovered an unexplored design space. The absence of precedent is itself information. The ontology provides the grammar; the retrieved examples provide the vocabulary. Together, they enable generation that speaks the language of game design fluently.
7  Ontology as Validation Contract
Generation is not enough. LLMs are notoriously agreeable: they produce output that looks valid without ensuring it is valid. The ontology serves not only as a generation schema but as a validation contract: a set of constraints that generated designs must satisfy to count as coherent.
7.1  The Yes-Man Problem
An LLM asked to generate a deck-building game may mention deck_building as a mechanism yet include no cards in the components, producing structurally incoherent output that sounds like a deck-building game without being one. Pydantic validation catches type errors and missing fields, but semantic coherence (the relationship between mechanisms and components) requires deeper validation.
7.2  Ontological Constraint Checking
We encode mechanism–component dependencies as validation functions. Listing 3 shows the core checker: a dictionary maps each mechanism to the component field it requires (deck building $\to$ cards, area control $\to$ board, worker placement $\to$ tokens). When a mechanism is present but its required component is absent, the validator flags the incoherence.
1def _check_mechanism_requirements(
2 self, design: GameOntology
3) -> List[str]:
4 issues = []
5 requirements = {
6 MechanismType.DECK_BUILDING:
7 ("card_types",
8 "Deck building needs cards"),
9 MechanismType.AREA_CONTROL:
10 ("board_description",
11 "Area control needs board"),
12 MechanismType.WORKER_PLACEMENT:
13 ("tokens",
14 "Workers need tokens"),
15 }
16 for mech in design.primary_mechanisms:
17 if mech in requirements:
18 field, msg = requirements[mech]
19 val = getattr(
20 design.components, field, None)
21 if not val:
22 issues.append(msg)
23 return issues
Listing 3: Ontology validator excerpt checking mechanism–component coherence.
7.3  Iterative Refinement
DSPy’s Assert mechanism enables inline validation: when dspy.Assert(is_valid, error_message) fails, DSPy automatically retries generation with the error message included as feedback. For complex validation requiring multiple attempts, we implement an explicit refinement loop: generate, validate, refine with specific issue descriptions, and repeat up to a configurable maximum. The validator returns both a boolean and a list of specific issues, enabling targeted correction rather than wholesale regeneration.
This architecture treats the ontology as a contract between the generation system and downstream consumers. Downstream systems (whether human designers reviewing output, automated balancing tools, or game engines instantiating the design) can rely on structural guarantees: a deck-building game will have cards; a cooperative game will not have direct conflict; the end condition will be specified.
8  Case Study: Mycelium: The Deep
Theory is persuasive; demonstration is convincing. We trace a complete generation through the Generative Ontology pipeline, from initial theme to playable game design.
8.1  Input
We provide a theme chosen to be unusual enough that retrieval finds no direct precedents, forcing the system to synthesize rather than imitate:
Theme: “Bioluminescent fungi competing for dominance in a deep cave ecosystem”
Constraints: 2–4 players, medium complexity, cooperative, 60–90 minutes
8.2  Phase 1: Mechanics Architect
The Mechanics Architect analyzes the theme and selects three core mechanisms: RESOURCE_MANAGEMENT (spore allocation across growth opportunities), AREA_CONTROL (territory expansion through the cave network), and ENGINE_BUILDING (specialized fungal structures providing ongoing benefits). The turn structure uses simultaneous planning (players place spore tokens secretly) followed by sequential resolution. The agency justification notes that players must balance short-term growth against long-term network health, with engine choices creating divergent strategies even in cooperative play.
8.3  Phase 2: Theme Weaver
The Theme Weaver integrates narrative, titling the game Mycelium: The Deep. The setting is an ancient limestone cave system where bioluminescent fungi are the only light source, slowly flooding as underground rivers shift. Thematic integration maps each mechanism to its fictional counterpart: spore placement is reproductive strategy; area control is securing nutrient pathways; engine building is developing specialized structures (fruiting bodies for points, rhizomorphs for fast movement, symbiotic bacteria for defense). The antagonist is the rising water table, creating cooperative tension without player-versus-player conflict.
8.4  Phase 3: Component Designer
The Component Designer specifies 30 modular hex cave tiles arranged vertically (deeper at the bottom), three card types (Event, Structure, Objective), example cards (“Flash Flood: Water rises 2 levels this round”; “Fruiting Body: Costs 3 growth, produces 1 VP per round”), and tokens including fungal growth cubes (15 per player in 4 colors), spore tokens (3 per player, regenerating each round), and flood markers (20 blue cubes advancing upward).
Figure 5: Case study: the design conversation. Each agent builds on the previous, with critics providing feedback that drives refinement before final assessment.
8.5  Phase 4: Balance Critic
The Balance Critic identifies three issues. First, a runaway leader problem in cooperative play: if one player’s strain grows faster, they dominate decisions (severity: moderate; recommendation: add a network synergy rule so connected strains share benefits). Second, arbitrary flood timing: random event cards determining flood speed leaves players with no agency over the core tension (severity: moderate; recommendation: make flood a predictable escalation of one level per round, with events as modifiers). Third, engine building as a potential trap: structures costing growth may cause builders to fall behind (severity: minor; recommendation: structures should provide an immediate small benefit plus an ongoing ability).
8.6  Phase 5: Refinement
The refinement agent addresses both moderate-severity issues: it adds a Network Synergy rule (adjacent strains share engine abilities), changes flood timing to a predictable one-level-per-round baseline with event modifiers, and rebalances structures to provide an immediate 1 VP plus their ongoing ability.
8.7  Phase 6: Fun Factor Judge
The Fun Factor Judge rates the refined design 7/10. Engagement hooks include persistent tension (“Will we reach the surface in time?”), simultaneous placement creating reveal anticipation, and structure building as achievement. Tension moments include the visible flood countdown, event card reveals, and decisions about saving endangered colonies versus abandoning them. Replayability stems from modular board setup, different strain combinations ($4!=24$ team configurations), and event deck variability.
8.8  Output Summary
Table 5 presents key fields of the generated design.
Table 5: Abbreviated output: Mycelium: The Deep.
Field Value
Title Mycelium: The Deep
Type Cooperative
Goal Grow fungal network from cave depths to surface before flood reaches highest colony
End Victory: growth cube reaches surface tile. Defeat: flood reaches highest cube row
Mechanisms Resource management, area control, engine building
Turn loop Place spores $\to$ Reveal $\to$ Grow $\to$ Event $\to$ Flood rises
Players 2–4, role-differentiated (Spreader, Rooted, Lumens, Toxic)
Components 30 hex tiles, 3 card types, 60 growth cubes, 12 spore tokens, 20 flood markers
8.9  Analysis
Without the ontology schema, an LLM generating “a cooperative game about bioluminescent fungi” would likely produce vague victory conditions (“escape the cave somehow”), inconsistent mechanisms (mentioning deck-building without specifying cards), missing components, unbalanced elements, and disconnected theme. Generative Ontology enforced complete goal specification via minimum-length constraints, coherent mechanism–component alignment via the validation function, thematic integration via the Theme Weaver’s explicit mapping, playability basics via required fields for turn structure and end condition, and balance review via the critic agent’s adversarial posture. The output is not merely plausible; it is playable. A designer could take this output, build a prototype, and begin playtesting immediately.
9  Generalizability
We developed Generative Ontology through tabletop games because games have rich ontological structure and connect to our previous work. But the pattern extends to any domain with coherent structure, where experts share vocabulary, where valid outputs satisfy definable constraints, and where knowledge accumulates in exemplars.
Figure 6: Beyond games: any domain with a “validity grammar” is a candidate for Generative Ontology.
9.1  Cross-Domain Application
Table 6 illustrates eight domains where Generative Ontology applies, each with its ontological components, generation outputs, and validation criteria.
Table 6: Cross-domain applicability of Generative Ontology.
Domain Ontology Elements Generation Output Validation Criteria
Tabletop Games Mechanisms, components, dynamics Novel game designs Playability, balance, coherence
Music Composition Scales, progressions, forms Original compositions Harmonic validity, structure
Architecture Spaces, materials, flows, codes Building designs Structural integrity, compliance
Culinary Arts Ingredients, techniques, flavors Novel recipes Flavor balance, feasibility
Narrative Characters, arcs, themes Stories and plots Narrative coherence, pacing
Software Design Patterns, components, interfaces System architectures Consistency, coupling, coverage
Legal Documents Clauses, parties, obligations Contracts Legal validity, completeness
Scientific Protocols Methods, materials, controls Experimental designs Reproducibility, validity
The methodology follows a consistent six-step pattern regardless of domain: (1) domain analysis, identifying expert concepts, relationships, and validity criteria; (2) ontology encoding, mapping concepts to Pydantic models with validators; (3) generation signatures, defining typed input/output contracts with expert personas; (4) retrieval augmentation, indexing exemplars by ontology structure; (5) validation loop with structural, semantic, and adversarial checks plus iterative refinement; and (6) multi-agent specialization, assigning focused expertise and concerns per ontology subdomain.
9.2  When to Use Generative Ontology
The pattern is most valuable when four conditions hold: (1) the domain has structure: experts can articulate what makes outputs valid; (2) exemplars exist: a corpus of successful examples enables RAG and grounds generation; (3) validity is checkable: constraints can be encoded as validation functions; and (4) generation is valuable: the effort of building the ontology is justified by repeated generation needs. Game design exemplifies all four criteria: rich structure, thousands of documented games [1], clear validity criteria, and ongoing need for novel designs.
10  Conclusion and Future Work
We began with a question: what happens when ontology learns to create? The answer is Generative Ontology, a synthesis combining the structural precision of knowledge representation with the creative power of large language models [13].
10.1  Key Insights
Four insights emerged from this work:
The Grammar Paradox.
Constraints enable creativity. A poet needs grammar not because grammar limits poetry but because grammar makes poetry possible. Our ontology schema, with its required fields, enum constraints, and validation rules, does not restrict the LLM’s creativity but provides the scaffolding within which meaningful creativity occurs.
Neither Alone Suffices.
Traditional ontology without generation is static, a map that cannot build cities. LLM generation without ontology is unconstrained, fluent outputs that collapse under scrutiny. The synthesis produces outputs that are both creative and valid, both novel and coherent.
The Whiteheadian Connection.
Generative Ontology offers a concrete instantiation of Whitehead’s metaphysics [16, 6]. The ontology provides eternal objects, abstract patterns existing as potentials. Generation produces actual occasions, concrete instantiations where form meets matter. Creativity is the movement from potentiality to actuality.
The Pattern Generalizes.
Although we developed Generative Ontology through games, the pattern applies wherever structured domains exist. Music, architecture, cooking, narrative, software design: each has its grammar, its ontology, its space of valid outputs.
10.2  Future Directions
This work opens several paths for exploration. Optimization: DSPy’s optimization capabilities [10] (BootstrapFewShot, MIPRO) can improve generation quality over time by learning from successful outputs; the ontology validation function we defined enables principled optimization via reward shaping. Interactive design: a human designer could collaborate with the system, providing high-level creative direction while the system handles structural completeness, with the ontology serving as a shared language between human and AI. Cross-domain transfer: ontologies developed for one domain may inform generation in related domains; game mechanisms and narrative structures share conceptual space, and the abstraction of Generative Ontology may enable cross-pollination. Ontology learning: we assumed the ontology is given, but AI systems could learn ontologies from corpora of successful examples; the induction of generative grammars from exemplars is a compelling research direction.
The grammar does not write the poem. But without grammar, there is no poem to write. Generative Ontology gives AI the grammar it needs to create meaningfully within any structured domain. The GameGrammar system is available at https://github.com/bennycheung/GameGrammarCLI.