### v2

# Generative Ontology: When Structured Knowledge Learns to Create

Abstract
Traditional ontologies excel at describing domain structure but cannot generate novel artifacts. Large language models generate fluently but produce outputs that lack structural validity, hallucinating mechanisms without components, goals without end conditions. We introduce Generative Ontology, a framework that synthesizes these complementary strengths: ontology provides the grammar; the LLM provides the creativity.
Generative Ontology encodes domain knowledge as executable Pydantic schemas that constrain LLM generation via DSPy signatures. A multi-agent pipeline assigns specialized roles to different ontology domains: a Mechanics Architect designs game systems, a Theme Weaver integrates narrative, a Balance Critic identifies exploits, each carrying a professional “anxiety” that prevents shallow, agreeable outputs. Retrieval-augmented generation grounds novel designs in precedents from existing exemplars, while iterative validation ensures coherence between mechanisms and components.
We demonstrate the framework through GameGrammar, a system for generating complete tabletop game designs, and present three empirical studies. An ablation study (120 designs, 4 conditions) shows that multi-agent specialization produces the largest creative quality gains (fun $d=1.12$, strategic depth $d=1.59$; both $p<.001$), while schema validation eliminates structural errors ($d=4.78$). A benchmark comparison against 20 published board games reveals structural parity but a bounded creative gap (fun $d=1.86$): generated designs score 7–8 while published games score 8–9, with tension/drama and social interaction already at parity. A test-retest reliability study (ICC analysis, 50 evaluations) validates the LLM-based evaluator, with 7 of 9 metrics achieving Good-to-Excellent reliability (ICC 0.836–0.989).
The pattern generalizes beyond games. Any domain with expert vocabulary, validity constraints, and accumulated exemplars is a candidate for Generative Ontology. Constraints do not limit creativity but enable it: just as grammar makes poetry possible, ontology makes structured generation possible.
1  Introduction
Ontologies describe what is. They define concepts, relationships, and constraints that codify expert understanding of a domain. In our earlier work on tabletop game ontology [5], we built a structured vocabulary that could decompose Catan into mechanisms (resource trading, modular board, dice-driven production), components (hex tiles, resource cards, settlements), and player dynamics (competitive, negotiation-heavy, variable player count). That ontology gave us a precise language for analysis.
But analysis is not creation.
Understanding the structure of sonnets does not make one a poet. Knowing the rules of chess does not generate new games. A traditional ontology, for all its precision, remains passive, a tool for describing what already exists, not for imagining what could be. We ask: what happens when we transform that passive vocabulary into an active grammar?
1.1  The Problem: Structural Hallucination
Large language models (LLMs) excel at fluent generation. Ask an LLM to “design a deck-building game set in a haunted mansion” and it will produce paragraphs of creative output: themes, mechanisms, components, and victory conditions flowing freely. Yet this fluency hides a deeper problem: LLMs generate without understanding structure.
Consider the output for such a prompt: “Players explore Ravenshollow Manor, collecting ghost cards and building their decks. The game features a unique ‘fear mechanic’ where players manage their sanity. Victory goes to whoever escapes with the most treasure…” This sounds plausible. But what cards exist in the starting deck? How do players acquire new cards? What triggers the end of the game? How does the “fear mechanic” interact with deck composition? The LLM has generated the appearance of a game design without the substance. Handed to a playtester, it would collapse into confusion. This is the hallucination problem applied to creative domains: just as LLMs hallucinate facts, they hallucinate structures, producing outputs that sound valid without being valid.
Figure 1: A sample Game Ontology expressed as an OntoUML class diagram. The Game root entity decomposes into Player, Mechanism, and Component aggregates, each with typed relationships (hasPlayer, hasMechanism, hasComponent) that mirror the domain structure a Generative Ontology must capture and enforce.
1.2  Complementary Limitations
Table 1 summarizes the impasse. Traditional ontology has structure without creativity; the LLM has creativity without structure. Neither alone solves the problem of generating novel, valid designs.
Table 1: Traditional ontology versus pure LLM generation.
Strength Weakness
Ontology Precise, structured, validated Cannot generate novel outputs
LLM Creative, fluent, abundant Unstructured, hallucinated
The insight of Generative Ontology is that these limitations are complementary. What ontology lacks, LLMs provide. What LLMs lack, ontology provides. The synthesis, structured generation guided by ontological schemas, produces outputs that are both creative and valid.
1.3  Paper Overview
We introduce Generative Ontology and develop it through the domain of tabletop game design. Section 2 reviews related work. Section 3 defines the framework and its philosophical grounding. Sections 48 develop the technical architecture: executable schemas, DSPy operationalization, multi-agent pipelines, retrieval augmentation, and validation contracts. Section 9 presents three empirical studies evaluating the framework. Section 10 illustrates the pipeline with a complete case study. Section 11 argues the pattern generalizes beyond games. Section 13 concludes with future directions.
2  Related Work
2.1  Ontologies and LLMs
Ontology development has a well-established methodology [15], but recent research at the intersection of ontologies and LLMs proceeds primarily in the opposite direction from our approach. DRAGON-AI uses retrieval-augmented LLM generation to construct biomedical ontologies from text corpora [18], and OLLM demonstrates end-to-end ontology learning with language models. These systems use LLMs to generate ontologies; we use ontologies to guide generation.
Mehenni and Zouaq [13] demonstrate ontology-constrained generation for clinical summaries, the closest methodological precedent to our work, though their goal is hallucination reduction in summarization rather than structured creativity. Cooper [7] provides foundational work on constrained decoding that underpins our structured output approach.
2.2  LLMs in Game Design
Gallotta et al. [9] survey LLM applications in games broadly, while Becker et al. [4] present Boardwalk, a framework for generating board game code from existing rules, generating implementations rather than designs. Maleki and Zhao [12] survey procedural content generation methods, noting emerging LLM integration but without ontological constraints.
2.3  LLM-as-Judge Evaluation
Recent work has established LLMs as evaluators of creative output [20]. Concerns about evaluator reliability motivate our test-retest analysis: Shankar et al. [16] argue that “who validates the validators” is a critical question for LLM-based evaluation pipelines. We address this by reporting intraclass correlation coefficients (ICC) for all evaluation metrics [11, 17], demonstrating that our evaluator is a reliable instrument rather than assuming so.
2.4  Gap
Our contribution fills a specific gap: using ontology as a generative grammar for structured creative output, operationalized through executable schemas, multi-agent pipelines, and retrieval augmentation, with rigorous empirical evaluation including ablation, benchmarking, and evaluator validation.
3  Generative Ontology: The Synthesis
If traditional ontology is a map, Generative Ontology is a map that knows how to build new cities.
Generative Ontology is the practice of encoding domain knowledge as executable schemas that constrain and guide AI generation, transforming static knowledge representation into a creative engine.
When we built our tabletop game ontology [5], we created a structured vocabulary: game types, mechanisms, components, player interactions. This vocabulary enabled analysis of games such as Catan and Dune: Imperium with precision. The insight of Generative Ontology is that the same structure enabling understanding can enable creation. The categories that constrain analysis become the scaffolding for generation.
3.1  The Grammar of Games
A poet does not experience grammar as a limitation. Grammar is not what prevents poetry; it is what makes poetry possible. Without the structure of syntax, semantics, and form, there would be no sonnets, no haiku, no free verse pushing against convention. The rules are not the enemy of creativity; they are its condition of possibility.
The same principle applies to game design. When we ask an unconstrained LLM to “design a board game about space exploration,” we receive verbose descriptions that sound plausible but collapse under scrutiny. The AI might describe a “resource management mechanic” without specifying what resources exist, how they flow, or what makes acquiring them interesting. Generative Ontology provides the grammar. When we encode the game ontology as a schema, we give the LLM the structural vocabulary to be creative coherently. The schema declares: every game must have a goal, an end condition, mechanisms that create player choices, components that instantiate those mechanisms. Within those constraints, infinitely many games are possible. Without them, no valid game emerges.
Figure 2: LLM potential plus ontology constraints yields valid design. Generative Ontology transforms a passive vocabulary for description into an active grammar for creation.
Table 2 illustrates the difference concretely.
Table 2: Generation without versus with Generative Ontology.
Without Ontology With Generative Ontology
“Players collect resources and build things” Goal: first to 10 VP via settlements
“There’s some kind of trading” Mechanism: resource trading with 4:1 bank or negotiated rates
“The game ends eventually” End Condition: when any player reaches 10 VP or development deck exhausts
Fluent but vague Structured and playable
3.2  Philosophical Grounding
For readers familiar with our earlier exploration of process philosophy for AI agent design [6], Generative Ontology has a natural grounding in Alfred North Whitehead’s metaphysics [19]. Whitehead distinguished between eternal objects (pure forms, patterns, and structures that exist as potentials) and actual occasions (the concrete events where eternal objects find expression). Our game ontology is a collection of eternal objects: the abstract patterns of worker placement, deck building, area control. These patterns have no reality until they are instantiated in actual games. Generative Ontology is the process of moving from eternal objects to actual occasions: the ontology provides the forms; generation produces the instances. Creativity, in Whitehead’s view, is precisely this novel synthesis of eternal objects into actual occasions that have never existed before [3].
4  From Ontology to Executable Schema
Philosophy illuminates the path; engineering builds the road. The key insight is surprisingly direct: ontology classes map naturally to schema definitions, and schema definitions become the structured outputs that constrain LLM generation.
4.1  The Transformation
Our tabletop game ontology [5, 8] defined four core concepts: Game (the container with goal and end condition), Mechanism (turn structure, actions, resolution, uncertainty), Component (board, cards, tokens), and Player (roles and interactions). Each of these ontology classes becomes a Pydantic BaseModel, a typed, validated structure that LLMs must conform to when generating output.
Listing 1 shows the abbreviated GameOntology schema. The MechanismType enumeration reflects ontological categories developed over decades of game design taxonomy [8]. By encoding these as an enum, we ensure the LLM can only reference mechanisms within our shared vocabulary. Every field in GameOntology corresponds to an ontology concept. The goal and end_condition are required strings with minimum lengths, preventing vague outputs. The nested ComponentSet and PlayerDynamics models enforce structure at every level.
1from enum import Enum
2from typing import List, Literal, Optional
3from pydantic import BaseModel, Field
4
5class MechanismType(str, Enum):
6 WORKER_PLACEMENT = "worker_placement"
7 ACTION_POINTS = "action_points"
8 DECK_BUILDING = "deck_building"
9 RESOURCE_MGMT = "resource_management"
10 AREA_CONTROL = "area_control"
11 ENGINE_BUILDING = "engine_building"
12 SET_COLLECTION = "set_collection"
13 HIDDEN_INFO = "hidden_information"
14
15class GameOntology(BaseModel):
16 title: str = Field(description="Name of the game")
17 theme: str = Field(description="Setting and narrative")
18 game_type: Literal["cooperative","competitive",
19 "semi-cooperative"]
20 goal: str = Field(min_length=20,
21 description="Victory condition")
22 end_condition: str = Field(min_length=10,
23 description="What triggers game end")
24 primary_mechanisms: List[MechanismType] = Field(
25 min_length=2, max_length=4)
26 turn_structure: str
27 uncertainty_source: str
28 components: ComponentSet # nested model
29 players: PlayerDynamics # nested model
30 setup: str
31 core_loop: str
32 strategic_depth: str
Listing 1: Abbreviated game ontology Pydantic schema. Each field corresponds to an ontology concept; enums constrain the LLM to recognized mechanisms.
4.2  Ontology-to-Schema Correspondence
Table 3 formalizes the mapping. The correspondence is not accidental: every ontology class finds expression in the schema, and every ontological relationship (Game has Mechanisms, Game uses Components) becomes a nested field or list. The schema is the ontology made executable.
Table 3: Ontology-to-schema correspondence.
Concept Schema Type Purpose
Game Types Literal[...] Valid modes
Mechanisms List[Mechanism] Mechanics
Components ComponentSet Physical parts
Goal / End str (required) Playability
Players PlayerDynamics Interactions
4.3  Why Pydantic?
The choice of Pydantic [2] is deliberate. Declaring GameOntology as a BaseModel provides four critical capabilities: (1) type validation, where the LLM’s output is automatically validated against the schema; (2) self-documentation, where each Field description becomes part of the prompt context; (3) nested structure, where hierarchical ontologies map naturally to nested models; and (4) enum constraints, where the LLM operates within the ontological vocabulary and cannot invent mechanisms outside the taxonomy.
5  Operationalizing with DSPy
We have the ontology encoded as Pydantic schemas. Now we need an engine to drive generation. DSPy (Declarative Self-improving Python) [10] treats LLM interactions as composable, typed operations rather than monolithic prompt templates. This philosophy aligns with Generative Ontology: DSPy signatures are ontological contracts.
5.1  Signatures as Ontological Contracts
A DSPy signature declares inputs, outputs, and a docstring that establishes the LLM’s persona. Listing 2 shows the core generation signature. The docstring becomes the system prompt; the input fields capture what the user provides; the output field declares that the result must conform to our GameOntology schema.
1import dspy
2from game_ontology import GameOntology
3
4class GenerativeOntologySignature(dspy.Signature):
5 """You are an expert tabletop game designer
6 with deep knowledge of mechanisms, player
7 dynamics, and component design. Design a
8 complete, playable game based on the theme.
9 Every field must be specific, not vague."""
10
11 theme = dspy.InputField(
12 desc="Thematic concept for the game")
13 design_constraints = dspy.InputField(
14 desc="Player count, complexity, time",
15 default="")
16 game_design: GameOntology = dspy.OutputField(
17 desc="Complete game design per schema")
Listing 2: DSPy generation signature. The typed output field binds generation to the ontological schema.
5.2  Execution Flow
When we execute this signature, DSPy performs the following sequence: (1) constructs a prompt incorporating the docstring, field descriptions, and the full Pydantic schema structure; (2) sends the prompt to the configured LLM; (3) parses the response into the GameOntology model; (4) validates that all required fields are present and correctly typed; and (5) returns the validated result or raises an error for retry.
DSPy’s ChainOfThought module extends this by instructing the model to reason through the design before committing to structured output. Encapsulating logic in a DSPy Module enables composition of multiple generation steps, post-generation validation, and retry logic. The module adds a layer of semantic validation beyond what Pydantic provides: the schema ensures structural correctness (all fields present, correct types), while the module ensures coherence: that mechanisms match components and that game type aligns with player dynamics.
6  Multi-Agent Ontology Pipeline
A single LLM call, even with our ontology schema, carries the entire burden of game design: it must simultaneously consider mechanisms, theme, components, balance, and player experience. Human design teams do not work this way. Generative Ontology enables the same division of labor by decomposing the ontology into domains and assigning specialized agents to each.
6.1  Agent Roster and Professional Anxieties
Table 4 lists the five agents. The “Anxiety” column is crucial: each agent is not merely assigned a domain but is given a concern, a professional worry that shapes its generation and critique. This anxiety-driven design prevents the “yes-man” tendency of LLMs to produce plausible but shallow output.
Table 4: Agent roster with ontology domains and professional anxieties.
Agent Domain Anxiety
Mechanics Architect Mechanisms, turn structure “Is there meaningful player agency?”
Theme Weaver Narrative, setting “Does the theme feel alive in every mechanism?”
Component Designer Cards, tokens, board “Can players manipulate this smoothly?”
Balance Critic Cross-domain analysis “What breaks when optimized?”
Fun Factor Judge Player experience “Would I want to play this again?”
Figure 3: Multi-agent pipeline architecture. Generative agents build sequentially (Mechanics $\to$ Theme $\to$ Components); critic agents evaluate in parallel; refinement addresses identified issues before final assessment.
6.2  Pipeline Architecture
The pipeline, illustrated in Figure 3, proceeds in four phases:
  1. 1.
    Sequential Generation. The Mechanics Architect produces core systems given the theme and constraints. The Theme Weaver receives these mechanics and integrates narrative. The Component Designer receives both and specifies physical elements.
  2. 2.
    Critical Evaluation. The Balance Critic examines the assembled design for exploits, dominant strategies, and interaction gaps.
  3. 3.
    Refinement. If moderate or severe issues are found, a refinement agent addresses specific recommendations while preserving the design’s strengths.
  4. 4.
    Experience Assessment. The Fun Factor Judge evaluates engagement hooks, tension moments, satisfaction sources, and replayability.
7  Retrieval-Augmented Generation
Our Generative Ontology can create novel game designs, but creativity without context risks reinventing solutions that already exist. The most interesting creative work happens in conversation with what has come before.
Figure 4: Knowledge System architecture. A unified Retriever Service dispatches queries to ChromaDB (semantic vector search) or SQLite (structured fallback), drawing on a BGG corpus of 1,767 games with 384-dimensional embeddings, mechanism types, and filterable metadata.
7.1  Two-Phase Retrieval
BoardGameGeek [1] contains data on over 100,000 tabletop games. We structure this corpus according to our ontology and employ a two-phase retrieval strategy: (1) ontology filtering, which narrows candidates to games sharing target mechanisms from the MechanismType taxonomy; then (2) semantic ranking, which ranks the filtered set by embedding similarity to the thematic query. This ensures retrieved games are both mechanically relevant and thematically resonant.
7.2  Grounded Creativity
RAG transforms Generative Ontology from invention to informed invention. The system knows what mechanisms have been successfully combined, how similar themes have been approached, and what design patterns work for given player counts and complexity levels. The ontology provides the grammar; the retrieved examples provide the vocabulary. Together, they enable generation that speaks the language of game design fluently.
8  Ontology as Validation Contract
Generation is not enough. The ontology serves not only as a generation schema but as a validation contract: a set of constraints that generated designs must satisfy to count as coherent.
8.1  Ontological Constraint Checking
We encode mechanism–component dependencies as validation functions. Listing 3 shows the core checker: a dictionary maps each mechanism to the component field it requires (deck building $\to$ cards, area control $\to$ board, worker placement $\to$ tokens). When a mechanism is present but its required component is absent, the validator flags the incoherence.
1def _check_mechanism_requirements(
2 self, design: GameOntology
3) -> List[str]:
4 issues = []
5 requirements = {
6 MechanismType.DECK_BUILDING:
7 ("card_types",
8 "Deck building needs cards"),
9 MechanismType.AREA_CONTROL:
10 ("board_description",
11 "Area control needs board"),
12 MechanismType.WORKER_PLACEMENT:
13 ("tokens",
14 "Workers need tokens"),
15 }
16 for mech in design.primary_mechanisms:
17 if mech in requirements:
18 field, msg = requirements[mech]
19 val = getattr(
20 design.components, field, None)
21 if not val:
22 issues.append(msg)
23 return issues
Listing 3: Ontology validator excerpt checking mechanism–component coherence.
8.2  Iterative Refinement
DSPy’s Assert mechanism enables inline validation: when dspy.Assert(is_valid, error_message) fails, DSPy automatically retries generation with the error message included as feedback. This architecture treats the ontology as a contract between the generation system and downstream consumers. Downstream systems (whether human designers, balancing tools, or game engines) can rely on structural guarantees: a deck-building game will have cards; a cooperative game will not have direct conflict; the end condition will be specified.
9  Empirical Evaluation
The preceding sections describe the Generative Ontology framework. We now evaluate it through three complementary studies: an ablation study isolating the contribution of each architectural component, a benchmark comparison against published board games, and a test-retest reliability analysis validating our evaluation instrument.
9.1  Experimental Setup
Generation Model. All designs were generated using Claude Sonnet 4 (model ID claude-sonnet-4-20250514, temperature 0.7) via DSPy.
Evaluation Pipeline. Each generated design is evaluated through two complementary lenses: (1) structural metrics, computed deterministically by checking the ontology for completeness (are all required fields present?), alignment (do mechanisms match components?), and consistency (are there contradictions?); and (2) creative metrics, scored by an LLM evaluator (“Design Coach”) on seven dimensions: strategic depth, tension/drama, player agency, replayability, social interaction, elegance, and thematic cohesion, each on a 1–10 scale. The Design Coach also provides an overall fun rating (1–10) and engagement variance (standard deviation across dimension scores). The evaluator uses the same model at temperature 0 for near-deterministic scoring.
Theme Prompts. Ten standardized theme prompts span a range of complexity levels and game types (see Appendix A). Each prompt specifies a theme, player count, complexity level, and target play time.
Experimental Conditions. Four conditions isolate the contribution of each pipeline component (Table 5).
Table 5: Experimental conditions for the ablation study.
Condition Configuration
C1 (Baseline) Raw LLM generation, no schema or ontology
C2 (Schema) Pydantic schema validation only
C3 (Ontology) Schema + ontology constraints, single agent
C4 (Pipeline) Full pipeline: schema + ontology + RAG + multi-agent + validation
9.2  Study 1: Ablation
Design. $4\text{ conditions}\times 10\text{ prompts}\times 3\text{ replications}=120$ designs, all evaluated through the full pipeline.
[…]
Table 6: Structural metrics by condition.
Condition Errors Compl.
C1 (Baseline) 5.03 100%
C2 (Schema) 0.10 100%
C3 (Ontology) 0.00 83%
C4 (Pipeline) 0.00 100%
Figure 5: Structural metrics by condition. Left: mean consistency errors drop from 5.03 (C1) to zero (C3, C4). Right: completeness is 100% for all conditions except C3 (83%), where single-agent generation misses detail fields.
[…]
Figure 6: Creative dimension score distributions by condition. C4 (Pipeline) shows the highest and tightest distributions on most dimensions. Replayability and social interaction do not discriminate between conditions.
[…]
Figure 7: Cost analysis. Left: generation tokens per design. Right: tokens per fun point. C4 costs ${\sim}8.5\times$ more but achieves the highest fun rating.
Summary. Structural validity is solved by schema validation. Creative quality arises from the interaction between constraint enforcement and architectural specialization: constraints alone suppress richness; specialization restores and enhances it.
9.3  Study 2: Benchmark Comparison
Design. We compare 30 C4-generated designs from Study 1 against 20 published board games evaluated through the same pipeline. The reference set spans three complexity tiers: light (Azul, Splendor, Love Letter, Kingdomino, Codenames), medium (Wingspan, Everdell, 7 Wonders, Concordia, Spirit Island), and heavy (Brass: Birmingham, Gaia Project, Twilight Imperium 4E, Through the Ages, Gloomhaven), plus five pre-existing ontologies (Catan, Root, Dune Imperium, Terraforming Mars, Ticket to Ride). See Appendix B for the full list. Each real game was reverse-engineered into a complete GameGrammar ontology and evaluated through the identical Design Coach pipeline.
Structural Parity. Both groups achieve 100% completeness. Generated designs have fewer consistency errors (1.27 vs. 2.80, $d=0.76$, $p<.01$). The higher error count for real games reflects richer component vocabularies (specialized tokens, standees, custom boards) that extend beyond the checker’s standard categories.
Creative Gap. Table 7 reports creative dimension scores. Published games score significantly higher on 5 of 7 dimensions, with the largest effects on elegance ($d=1.72$) and fun ($d=1.86$). Crucially, tension/drama ($d=0.35$, ns) and social interaction ($d=0.19$, ns) show no significant differences, suggesting the pipeline already handles these dimensions comparably.
Table 7: Creative dimensions: real vs. generated designs.
Dimension Real Gen. $d$ Sig.
Strategic Depth 8.85 8.13 0.94 **
Tension & Drama 8.50 8.20 0.35 ns
Player Agency 9.00 8.23 1.21 ***
Replayability 9.05 7.63 1.59 ***
Social Interaction 7.20 6.93 0.19 ns
Elegance 9.25 7.97 1.72 ***
Fun Rating 8.90 8.07 1.86 ***
$d$ = Cohen’s $d$; ns = not significant; ** $p<.01$; *** $p<.001$.
Figure 8: Violin plots showing score distributions for all creative dimensions. Real games (red) cluster in the 8–9 range while generated designs (green) spread across 7–9. Tension/drama and social interaction show the most overlap.
Mechanism Similarity. Generated designs use a similar number of mechanisms (6.7 vs. 6.1, d=0.60d=-0.60) but fewer component types (7.8 vs. 11.6, $d=1.12$, $p<.01$). The pipeline generates designs within a mid-complexity band but does not yet produce the extremes: neither ultra-light party games nor epic-scale strategy games (Figure 9).
Figure 9: Mechanism count distributions. Both groups center around 6–7 mechanisms; generated designs show slightly more spread.
Summary. Generated designs achieve structural parity with published games but exhibit a bounded creative gap of approximately one point on a 10-point scale. This gap likely reflects properties that emerge from iterative playtesting (elegance and replayability) rather than design specification alone.
9.4  Study 3: Evaluator Reliability
An empirical evaluation is only as credible as its evaluation instrument. We assess the Design Coach’s test-retest reliability using intraclass correlation coefficients (ICC).
Design. Ten designs spanning the quality range (fun ratings 6–9, all four conditions) were each evaluated five times through the Design Coach, yielding 50 evaluations. We compute ICC(2,1), two-way random effects with single rater and absolute agreement [17], for each of the nine LLM-scored metrics.
Results. Table 8 reports ICC values. Seven of nine metrics achieve Good-to-Excellent reliability (ICC $\geq 0.75$). Four metrics rate Excellent ($\geq 0.90$): tension & drama (0.989), social interaction (0.983), thematic cohesion (0.975), and elegance (0.934). Three rate Good: fun rating (0.873), strategic depth (0.846), and player agency (0.836). Two metrics, replayability (0.578) and engagement variance (0.598), show moderate reliability.
Table 8: Test-retest reliability: ICC(2,1) per metric.
Metric ICC 95% CI Level
Tension & Drama .989 [.970, 1.00] Excellent
Social Interaction .983 [.960, .990] Excellent
Thematic Cohesion .975 [.940, .990] Excellent
Elegance .934 [.850, .980] Excellent
Fun Rating .873 [.730, .960] Good
Strategic Depth .846 [.680, .950] Good
Player Agency .836 [.660, .950] Good
Engagement Var. .598 [.330, .850] Moderate
Replayability .578 [.300, .840] Moderate
Figure 10: ICC values per metric with 95% confidence intervals. Green = Excellent ($\geq.90$), blue = Good (.75–.90), orange = Moderate (.50–.75).
Cross-Study Validation. The ICC results retroactively validate prior findings. Fun rating (ICC = 0.873) confirms that the fun-based comparisons in Studies 1 and 2 rest on stable measurements. The high ICC for tension & drama (0.989) and social interaction (0.983) confirms that the non-significant differences in Study 2 reflect genuine equivalence, not measurement noise. Conversely, replayability’s moderate ICC (0.578) suggests that its non-significance in Study 1 may partly reflect evaluator inconsistency, a property arguably better assessed through playtesting than design specification analysis.
Summary. The Design Coach is a reliable evaluation instrument. Score differences between designs reflect genuine quality differences, not evaluator drift.
10  Case Study: Mycelium: The Deep
To illustrate the pipeline concretely, we trace a single generation from theme to playable design. The input is a theme chosen to be unusual enough that retrieval finds no direct precedents:
Theme: “Bioluminescent fungi competing for dominance in a deep cave ecosystem”
Constraints: 2–4 players, medium complexity, cooperative, 60–90 minutes
The Mechanics Architect selects three core mechanisms: resource management (spore allocation), area control (cave network expansion), and engine building (specialized fungal structures). The Theme Weaver titles the game Mycelium: The Deep and maps each mechanism to its fictional counterpart: spore placement as reproductive strategy, area control as securing nutrient pathways, engine building as developing specialized structures. The Component Designer specifies 30 modular hex cave tiles, three card types, and tokens including fungal growth cubes, spore tokens, and flood markers. The Balance Critic identifies a runaway-leader problem and arbitrary flood timing; the refinement agent adds a Network Synergy rule and predictable flood escalation. The Fun Factor Judge rates the refined design 7/10. Table 9 summarizes the output.
Table 9: Abbreviated output: Mycelium: The Deep.
Field Value
Title Mycelium: The Deep
Type Cooperative
Goal Grow fungal network from cave depths to surface before flood reaches highest colony
End Victory: growth cube reaches surface tile. Defeat: flood reaches highest cube row
Mechanisms Resource management, area control, engine building
Turn loop Place spores $\to$ Reveal $\to$ Grow $\to$ Event $\to$ Flood rises
Players 2–4, role-differentiated
Components 30 hex tiles, 3 card types, 60 growth cubes, 12 spore tokens, 20 flood markers
Without the ontology schema, an LLM generating “a cooperative game about bioluminescent fungi” would likely produce vague victory conditions, inconsistent mechanism–component mappings, and disconnected theme. Generative Ontology enforced complete goal specification, coherent mechanism–component alignment, and thematic integration. The output is not merely plausible; it is playable.
11  Generalizability
We developed Generative Ontology through tabletop games because games have rich ontological structure and connect to our previous work. But the pattern extends to any domain with coherent structure, where experts share vocabulary, where valid outputs satisfy definable constraints, and where knowledge accumulates in exemplars.
Figure 11: Beyond games: any domain with a “validity grammar” is a candidate for Generative Ontology.
Table 10 illustrates eight domains where Generative Ontology applies. The pattern is most valuable when four conditions hold: (1) the domain has structure: experts can articulate what makes outputs valid; (2) exemplars exist: a corpus of successful examples enables RAG; (3) validity is checkable: constraints can be encoded as validation functions; and (4) generation is valuable: the effort of building the ontology is justified by repeated generation needs.
Table 10: Cross-domain applicability of Generative Ontology.
Domain Ontology Elements Generation Output Validation Criteria
Tabletop Games Mechanisms, components, dynamics Novel game designs Playability, balance, coherence
Music Composition Scales, progressions, forms Original compositions Harmonic validity, structure
Architecture Spaces, materials, flows, codes Building designs Structural integrity, compliance
Culinary Arts Ingredients, techniques, flavors Novel recipes Flavor balance, feasibility
Narrative Characters, arcs, themes Stories and plots Narrative coherence, pacing
Software Design Patterns, components, interfaces System architectures Consistency, coupling, coverage
Legal Documents Clauses, parties, obligations Contracts Legal validity, completeness
Scientific Protocols Methods, materials, controls Experimental designs Reproducibility, validity
12  Discussion
12.1  The Constraint Paradox in Data Form
Our ablation results provide empirical evidence for what we term the constraint paradox: adding constraints (C2, C3) eliminates structural errors without improving creative quality, yet the combination of constraints with architectural specialization (C4) produces the largest creative gains. Constraints alone suppress richness; constraints plus specialization enable it. This suggests a Constraint–Architecture Interaction Model: creative quality is a function of constraint expressiveness multiplied by architectural specialization, not simply additive.
12.2  The Gap as Feature
The one-point creative gap between generated and published designs (Study 2) is, paradoxically, encouraging. Published games represent decades of iterative playtesting, community feedback, and designer refinement, processes that no single generation pass can replicate. That the pipeline produces designs in the 7–8 range (“good, playable, interesting”) rather than the 8–9 range (“polished, elegant, replayable”) suggests the gap reflects not a fundamental limitation of the approach but the absence of iterative refinement. The dimensions where parity is already achieved (tension/drama and social interaction) may be those most amenable to specification-time design, while the gap dimensions (elegance and replayability) may require iterative playtesting to optimize.
12.3  The Whiteheadian Connection
Generative Ontology offers a concrete instantiation of Whitehead’s metaphysics [19, 6]. The ontology provides eternal objects, abstract patterns existing as potentials. Generation produces actual occasions, concrete instantiations where form meets matter. Creativity is the movement from potentiality to actuality. Our empirical results show this movement is not random: the constraints shape the space of possible actualizations, and architectural specialization navigates that space effectively.
12.4  Limitations
Several limitations qualify our findings. LLM-as-evaluator: While our ICC analysis demonstrates evaluator reliability, it does not establish validity: LLM scores may be consistently biased. Human calibration (comparing LLM scores to expert judgment) remains future work. No playtesting: Our evaluation assesses design specifications, not gameplay. Properties like elegance and replayability may manifest differently in play than in specification. Reference set: The 20 benchmark games, while spanning three complexity tiers, represent a convenience sample of well-known titles. A larger, more diverse reference set would strengthen generalizability. Single model: All generation and evaluation used Claude Sonnet 4. Cross-model evaluation would assess robustness. Temperature sensitivity: Evaluator ICC was measured at temperature 0; reliability at higher temperatures is untested.
13  Conclusion and Future Work
We began with a question: what happens when ontology learns to create? The answer is Generative Ontology, a synthesis combining the structural precision of knowledge representation with the creative power of large language models [14].
Four insights emerged from this work:
The Grammar Paradox. Constraints enable creativity. Our ablation study confirms this empirically: schema validation eliminates structural errors ($d=4.78$) while multi-agent specialization produces the largest creative gains (fun $d=1.12$, depth $d=1.59$).
Neither Alone Suffices. Traditional ontology without generation is static. LLM generation without ontology is unconstrained. The benchmark comparison shows that the synthesis produces designs that are structurally equivalent to published games, with a bounded creative gap that likely reflects the absence of iterative playtesting rather than fundamental architectural limitations.
Reliable Measurement Matters. Our test-retest analysis (ICC 0.578–0.989, median 0.873) demonstrates that LLM-based evaluation can be reliable when properly constrained. The cross-study validation differentiates “non-significant due to noise” from “non-significant due to genuine equivalence,” a methodological contribution applicable beyond game design.
The Pattern Generalizes. Although we developed Generative Ontology through games, the pattern applies wherever structured domains exist. Music, architecture, cooking, narrative, software design: each has its grammar, its ontology, its space of valid outputs.
Future directions include human calibration (validating LLM scores against expert judgment), iterative design loops (multiple generation–evaluation–refinement cycles to close the creative gap), cross-model evaluation (assessing robustness across LLMs), and DSPy optimization (using the ontology validation function as a reward signal for automatic prompt tuning).
The grammar does not write the poem. But without grammar, there is no poem to write. Generative Ontology gives AI the grammar it needs to create meaningfully within any structured domain. The GameGrammar system is available at https://github.com/bennycheung/GameGrammarCLI.
Appendix A Theme Prompts
Table 11 lists the ten standardized theme prompts used in the ablation study. Each prompt was run under all four experimental conditions with three replications, yielding 120 total designs.
Table 11: Standardized theme prompts for the ablation study.
# Constraints Theme
1 2–4p, competitive, medium, 45–60min Bioluminescent fungi competing in a cave ecosystem
2 2–4p, competitive, med-heavy, 60–90min Victorian-era detective solving mysteries through deduction
3 3–5p, competitive, med-heavy, 60–90min Space traders negotiating routes between planets
4 2–4p, competitive, light, 30–45min Medieval knights jousting in a tournament
5 2–4p, cooperative, medium, 45–60min Deep sea divers recovering treasure from a sunken ship
6 2–4p, competitive, medium, 45–60min Rival chefs competing in a cooking competition
7 2–5p, competitive, med-heavy, 90–120min Ancient civilizations building wonders of the world
8 2–4p, cooperative, medium, 30–45min Escape room where the walls are closing in
9 2–4p, competitive, medium, 45–60min Fairy tale creatures vying for control of an enchanted forest
10 2–4p, cooperative, heavy, 90–120min Time travelers preventing paradoxes across historical eras
Appendix B Reference Games
Table 12 lists the 20 published board games used as the reference set in Study 2. Games were selected to span three complexity tiers (light, medium, heavy) and represent widely recognized titles with established design reputations. Each game was reverse-engineered into a complete GameGrammar ontology (25 required fields plus 3 balance parameters) and evaluated through the identical Design Coach pipeline used for generated designs.
Table 12: Reference games for the benchmark study (Study 2).
# Game Designer(s) Tier
1 Azul Michael Kiesling Light
2 Codenames Vlaada Chvátil Light
3 Kingdomino Bruno Cathala Light
4 Love Letter Seiji Kanai Light
5 Splendor Marc André Light
6 7 Wonders Antoine Bauza Medium
7 Catan Klaus Teuber Medium
8 Concordia Mac Gerdts Medium
9 Dune: Imperium Paul Dennen Medium
10 Everdell James A. Wilson Medium
11 Root Cole Wehrle Medium
12 Spirit Island R. Eric Reuss Medium
13 Terraforming Mars Jacob Fryxelius Medium
14 Ticket to Ride Alan R. Moon Medium
15 Wingspan Elizabeth Hargrave Medium
16 Brass: Birmingham Gavan Brown & Matt Tolman Heavy
17 Gaia Project Jens Drögemüller & Helge Ostertag Heavy
18 Gloomhaven Isaac Childres Heavy
19 Through the Ages Vlácil Chvátil Heavy
20 Twilight Imperium 4E Dane Beltrami & Corey Konieczka Heavy